What is the enthalpy change when 5.00 g of C3H6 is burned according to the equation: 2C3H6 + 9O2 --> 6CO2 +6H2O + 4116.0 kJ
1
Expert's answer
2020-04-18T07:06:56-0400

M(C3H6) = 12*3 + 1*6 = 42 g/mol

n(C3H6) = m/M = 5g/42g/mol = 0.119 mol

According to the equation

2 mol C3H6 - 4116 kJ

0.119 mol - x

Enthalpy change = 0.119*4116/2 = 244.9 kJ